OpenStudy (anonymous):

integrate sinx/sqrt(1+cosx) dx

OpenStudy (anonymous):

Use the substitution \[u=1+\cos x\] Then \[du = -\sin x\, dx\] and so the integral becomes \[-\int\limits_{}^{} u^{-1/2} du\] which can be solved easilty to give \[-2u^{1/2} + c\] and so the original integral is \[-2\sqrt{1+\cos x} + c\]